OpenStudy (anonymous):

HELP PLEASE! If f:R^2->R^3 and g:R^3->R and h = g◦f, h(uv)=g(f(u,v)) for every (u,v) E R^2. How do you calculate derivative h/u in point (a,b)?

OpenStudy (mathmate):

\[Given:\ f: R^2 \rightarrow R^3\ and \ g:R^3 \rightarrow R\]and \[h = g◦f,\ h(u,v)=g(f(u,v))\ \forall (u,v)\in R^2\]( please confirm correction of typo in question from "h(uv)" to "h(u,v)" ) Assuming f, g and h are continuous over R, then we can use the chain rule to find the derivative of h: \[\frac{dh}{du} = \frac{dg}{df}\frac{df}{du}\] and evaluate the derivative at (a,b).

OpenStudy (anonymous):

yes, i made a mistake of typo. so this mean in point (a,b) should be: